# The Inhibitor: ReLU and Addition-Based Attention for Efficient Transformers

### v1

@@ Abstract
To enhance the computational efficiency of quantized Transformers, we replace the dot-product and Softmax-based attention with an alternative mechanism involving addition and ReLU activation only. This side-steps the expansion to double precision often required by matrix multiplication and avoids costly Softmax evaluations but maintains much of the core functionality of conventional dot-product attention. It can enable more efficient execution and support larger quantized Transformer models on resource-constrained hardware or alternative arithmetic systems like homomorphic encryption. Training experiments on four common benchmark tasks show test set prediction scores comparable to those of conventional Transformers with dot-product attention. Our scaling experiments also suggest significant computational savings, both in plaintext and under encryption. In particular, we believe that the ReLU and addition-based attention mechanism introduced in this paper may enable privacy-preserving AI applications operating under homomorphic encryption by avoiding the costly multiplication of encrypted variables.
1 Introduction
@@
Transformer models [1] have achieved great success in various machine-learning tasks and are the foundation for modern large-scale pre-trained Language Models such as BERT [2], RoBERTa [3], XLNet [4], Transformer-XL [5], and the GPT family [6, 7]. These models have enabled leaps in performance on many NLP tasks, as for example experienced by the millions that use the ChatGPT application [8]. Similarly, the Vision-Transformer (ViT) [9], and similar models outperform CNNs on many computer vision tasks such as image classification [10] and object detection [11].
@@
The impressive gains in performance, however, come at a price. State-of-the-art Transformers are extremely large, having parameters in the hundreds of billions. The requirements for memory and computational power have become prohibitive, not only for deployment on resource-constrained embedded systems, but sometimes even in data centers, and their considerable energy usage is widely discussed in society. Neural network quantization [12, 13, 14] is a method to address this problem, as it reduces memory usage by employing lower-bit precision for weight and activation tensors. This not only decreases inference time but also increases energy efficiency through the adoption of low-bit fixed-point arithmetic instead of floating-point operations [15]. Other approaches to improve Transformer’s scalability and environmental sustainability include Pruning [16, 17], Knowledge Distillation [18, 19] and Neural Architecture Search [20]. In this work, we propose changes to the architecture’s fundamental building blocks to take steps towards the same end.
@@
Transformers use an attention mechanism to identify connections between different elements in a sequence, which conventionally takes the dot-product between the query and key matrices before passing the results through a Softmax function. Many alternative formulations have been proposed, such as the Fastformer [21] based on additive attention, or the ReLUformer [22] that uses ReLU activation in place of Softmax. These architectures share elements with what is proposed in this work, however, they still apply attention by a matrix multiplication with the value matrix. Extending our earlier work with ReLU and addition-based gated RNNs [23], we here propose the Inhibitor attention mechanism, which is designed to not rely on variable-to-variable multiplication and Softmax activation. These operations are more costly than constant-to-variable multiplications (also known as literal multiplication) and ReLU on many computing hardware. Even more so, they are particularly challenging operations under homomorphic encryption, an arithmetic system that permits calculation with encrypted variables without access to the secret key [24].
2 Method
@@
which is then used to form a weighted sum of values as output, $H=S\,V$, exploiting that each row of $S$ is normalized to sum to one. The FFN has two fully connected layers with weight matrix multiplication and ReLU activation,
@@
where $W_{1},W_{2}$ are weight matrices and $b_{1},b_{2}$ are bias vectors. We use the short-hand notation $x^{+}=\max(0,x)$ for the ReLU function. Each Transformer block typically also has one or more layer normalizations [25] that stabilize the gradient flow and contribute to regularizing the training.
The Inhibitor.
We propose to alternatively calculate attention scores as
@@ 3 Results
The adding problem was introduced by [26] to test long-term memory for RNNs. It takes two inputs: a sequence of random numbers and a two-hot sequence (which, in our experiments, each is 100 long). The ground truth is simply the dot-product of the two inputs as vectors, which, rather obviously, is not a challenging task for a conventional (dot-prod-based) Transformer. It is encouraging that the addition-based Inhibitor performs just as well on this task (which is hard for, e.g., simple RNNs to solve).
We also train one-layer Transformers for the MNIST handwritten digit recognition task [27] and IMDB movie-review sentiment analysis task [28], which are simple go-to benchmark task for image classification and text analysis, respectively. Although far from SotA, we note that our simple models achieve decent accuracy for both attention mechanisms, where the differences in performance are not significant.
The final task, labeled IAMW in Table 1, uses the IAM Handwriting Database [29], which is a collection of handwritten words written by more than 700 writers. The model first applies two convolutional layers followed by a Transformer and uses Connectionist Temporal Classification (CTC) loss [30] as an endpoint layer to predict the text. Edit distance [31] was used as the evaluation metric. Again, the differences are small.
@@
For the plaintext experiment, we used integer 16-bit arithmetics implemented in the Rust programming language, which gives detailed low-level control over circuits and supports advanced time benchmarking through the Criterion package. The encrypted implementation uses the TFHE scheme [32] for Fully Homomorphic Encryption on the Torus as supported in the Concrete library for Python [33]. TFHE permits the evaluation of arbitrary single-variate activation functions under homomorphic encryption through the Programmable Bootstrap [34]. We here used considerably smaller networks where the embedding dimension was limited to size 2. The encrypted circuits were implemented for integers with up to 8-bit precision.
@@
timing_short_tfhe.txt
@@
The results indicate that the proposed Inhibitor mechanism can have a significant advantage, with i) 30%–50% saving for the plaintext implementation on CPU as per Table 2, and ii) a factor 3–6 under encryption with TFHE as per table 3. Timing estimates are averaged over 20 repeated experiments and are significant at the 95%
@@ Conclusions.
The proposed Inhibitor mechanism allows straightforward quantization as equations 3 and 4 naturally support integer implementation. Our scaling experiments indicate that it has considerable potential for computational savings both for Transformers processing plaintext data as well as ciphertext.
While experiment results are promising on simple training tasks, for future work, it is necessary to examine performance under more challenging settings, for example, by pre-training a much larger Transformer model and testing on modern NLP and image recognition benchmark tasks.
@@
3.1 Signed Inhibitor
The inhibition mechanism described in equation 4 can only produce a non-negative attended value matrix $H$, while the conventional dot-product attention can also select from $V$ that can have both positive and negative entries. We can rewrite inhibition to handle signed values simply as
@@ 3.1 Signed Inhibitor
3.2 Implemention
A naive implementation of equation 3 and 4 relies on expanding the dimension of the broadcasted differences before passing the result through an $\mathrm{abs}$ or $\mathrm{ReLU}$ function and then summing over an inner dimension. This uses large amounts of memory and should be avoided.
@@ 3.2 Implemention
For the inhibition step of equation 4, we use a simple trick based on

### v2

@@ Abstract
To enhance the computational efficiency of quantized Transformers, we replace the dot-product and Softmax-based attention with an alternative mechanism involving addition and ReLU activation only. This side-steps the expansion to double precision often required by matrix multiplication and avoids costly Softmax evaluations but maintains much of the core functionality of conventional dot-product attention. It can enable more efficient execution and support larger quantized Transformer models on resource-constrained hardware or alternative arithmetic systems like homomorphic encryption. Training experiments on four common benchmark tasks show test set prediction scores comparable to those of conventional Transformers with dot-product attention. Our scaling experiments also suggest significant computational savings, both in plaintext and under encryption. In particular, we believe that the ReLU and addition-based attention mechanism examined in this paper may enable privacy-preserving AI applications operating under homomorphic encryption by avoiding the costly multiplication of encrypted variables.
Introduction
@@
Transformer models (Vaswani et al. 2017) have achieved great success in various machine-learning tasks and are the foundation for modern large-scale pre-trained Language Models such as BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), XLNet (Yang et al. 2019), Transformer-XL (Dai et al. 2019), and the GPT family (Radford et al. 2019; Brown et al. 2020). These models have enabled leaps in performance on many NLP tasks, as for example experienced by the millions that use the ChatGPT application (Schulman et al. 2022). Similarly, the Vision-Transformer (ViT) (Dosovitskiy et al. 2020), and similar models outperform CNNs on many computer vision tasks such as image classification (Liu et al. 2021) and object detection (Carion et al. 2020).
@@
The impressive gains in performance, however, come at a price. State-of-the-art Transformers are extremely large, having parameters in the hundreds of billions. The requirements for memory and computational power have become prohibitive, not only for deployment on resource-constrained embedded systems, but sometimes even in data centers, and their considerable energy usage is widely discussed in society. Neural network quantization (Zhou et al. 2016; Jacob et al. 2018; Krishnamoorthi 2018) is a method to address this problem, as it reduces memory usage by employing lower-bit precision for weight and activation tensors. This not only decreases inference time but also increases energy efficiency through the adoption of low-bit fixed-point arithmetics instead of floating-point operations (Horowitz 2014). Other approaches to improve Transformer’s scalability and environmental sustainability include Pruning (Han et al. 2015; Han, Mao, and Dally 2016), Knowledge Distillation (Hinton, Vinyals, and Dean 2015; Sanh et al. 2019) and Neural Architecture Search (Chen et al. 2021). This work discusses changes to the architecture’s fundamental building blocks towards the same end.
@@
Transformers use an attention mechanism to identify connections between different elements in a sequence, which conventionally takes the dot-product between the query and key matrices before passing the results through a Softmax function. Many alternative formulations have been proposed, such as the Fastformer (Wu et al. 2021) based on additive attention, or the ReLUformer (Shen et al. 2023) that uses ReLU activation in place of Softmax. These architectures share elements with the alternative formulation that is discussed in this work, however, they still apply attention by a matrix multiplication with the value matrix.
Extending our work with ReLU and addition-based gated RNNs (Brännvall et al. 2023), we proposed the Inhibitor attention mechanism (Brännvall 2024), which is designed to not rely on variable-to-variable multiplication and Softmax activation. These operations are more costly than constant-to-variable multiplications (also known as literal multiplication) and ReLU on many computing hardware. Even more so, they are particularly challenging operations under Fully Homomorphic Encryption (FHE), an arithmetic system that permits calculation with encrypted variables without access to the secret key. This article includes our previous results and gives more detail on the FHE implementation.
Preliminaries
TFHE, or Fully Homomorphic Encryption on the Torus (Chillotti et al. 2019), is distinguished from other FHE schemes because it proposes the so-called Programmable Bootstrapping (PBS) mechanism. The PBS does a blind rotation over a predefined table of all the permittable values in a message space, e.g [N,N1][-N,N-1], landing on the coefficient whose index is closest to the original cipher text $x$. By filling the lookup table with values of a function evaluated over the original message space, $f(x)$, one instead obtains the PBS, which both reduces noise and evaluates a univariate function on a discrete set of points. It is also relatively fast compared to alternative FHE schemes.
TFHE does not natively have multiplication between ciphertexts, but it can be constructed by the application of two PBS operations
$a\,b=\mathrm{PBS}\left(f;a+b\right)-\mathrm{PBS}\left(f;a-b\right)$ (1)
where by $\mathrm{PBS}\left(f;x\right)$, we mean applying a table that corresponds to the function $f$ with argument $x$, which is
$f(x)=\frac{x^{2}}{4}$ (2)
in this case of multiplication by PBS.
In order for the FHE scheme to be able to evaluate a circuit correctly, all the possible values that can occur for each component must be inside the allotted message space. The TFHE scheme operates on several different cipher-texts, namely LWE (Regev 2009), RLWE (Lyubashevsky, Peikert, and Regev 2013), GLWE (Brakerski, Gentry, and Vaikuntanathan 2012), and generalizations of GSW (Gentry, Sahai, and Waters 2013), each governed by specific parameters. For LWE ciphertexts, a dimension parameter is required, while GLWE ciphertexts rely on both polynomial size and dimension parameters.
A framework for parameter optimization in TFHE proposed by (Bergerat et al. 2023) distinguishes such macro-parameters from the micro-parameters used exclusively within FHE operators, such as the decomposition base and the number of decomposition levels required for a PBS. To optimize FHE operators, both noise and cost models come into play. The noise model simulates noise evolution across an operator, while the cost model serves as a metric to minimize encompassing factors like execution time, power consumption, or price. As the selection of micro and macro parameters significantly impacts cost and noise growth for each operation, they demand prudent and deliberate choices. For all numerical experiments in this paper, we use parameters automatically selected by the Concrete Python Compiler (Zama 2022) based on the TFHE parameter optimization framework by (Bergerat et al. 2023).
Privacy Preserving Machine Learning.
Despite the computational challenges, variants of Fully Homomorphic Encryption have been successfully applied to achieve Privacy Preserving Machine Learning (PPML). Such efforts include hyperplane decision-based classification, Naive Bayes classification, decision tree classification (Bost et al. 2015; Sun et al. 2020), classification and regression using Random Forests as demonstrated in Cryptotrees (Huynh 2020), and more recent developments of efficient nearest neighbor classifiers (Chakraborty and Zuber 2022).
Early approaches for Neural Networks like Cryptonets (Dowlin et al. 2016) focused on achieving blind, noninteractive classification by applying a leveled homomorphic encryption scheme (Brakerski, Gentry, and Vaikuntanathan 2012) to the network inputs and used the square function as an activation function replacement to accommodate the limitations of the underlying encryption scheme. Subsequent works (Zhang, Yang, and Chen 2016) adopted higher-degree polynomials that approximate conventional activation functions like sigmoid and tanh on a range. However, leveled homomorphic approaches have limited scalability, and as the number of layers in a neural network increases, the overall performance of homomorphic classification becomes prohibitive.
To overcome the limitations of previous approaches, a hybrid cryptographic scheme (Juvekar, Vaikuntanathan, and Chandrakasan 2018) was proposed that combined homomorphic encryption (HE) for processing linear layers with garbled circuits (GC) for computing activations in convolutional neural networks (CNNs). The garbled circuit is a cryptographic protocol that enables secure computation in which two parties can jointly evaluate a function over their private inputs (Yao 1982). This hybrid approach suffered from long inference latency due to slow GC-based activations, especially for tanh and sigmoid that required large circuits.
The emerging TFHE scheme (Chillotti et al. 2016) showed promise for noninteractive PPML because of its fast bootstrap procedure and capacity to evaluate look-up tables. Based on this scheme, FHE–DiNN, a novel framework for homomorphic evaluation of discretized neural networks, was proposed (Bourse et al. 2018). It featured linear complexity in the network depth and overcomes the scale limitations of leveled approaches, and applies the sign function for neuron activation by using bootstrapping. Empirical results demonstrate accurate classification by fully connected neural networks of encrypted MNIST dataset images with over 96% accuracy in under 1.7 seconds. Also using TFHE, a shift-accumulation-based convolutional deep neural network (SHE) (Lou and Jiang 2019) was proposed to overcome the limitations of using polynomial approximations by implementing ReLU activations and max pooling. It employed logarithmic quantization to replace expensive FHE multiplications with cheaper shifts and claim state-of-the-art inference accuracy and reduced inference latency on MNIST and CIFAR-10 image classification data sets.
Computational Efficiency.
Different operations within neural networks have distinct power and energy requirements, which can vary based on the underlying computer architecture (Parhami 2010). Addition is a relatively straightforward operation that can be executed in a single instruction. It is considered relatively inexpensive also under FHE.
Literal constants are fixed values that can be encoded in program instructions, while variables are undetermined at compile time. Literal multiplication can be optimized by the compiler using techniques like precomputed values or bit-shifting, resulting in faster execution and lower cost compared to variable multiplication also in clear text. Variable multiplication entails additional memory access to retrieve values, which can significantly impact energy consumption, as RAM access can be several orders of magnitude more energy-intensive than computation (Horowitz 2014). While multiplication by literals is natively supported in TFHE, the multiplication of two encrypted variables is constructed using two Programmable Bootstrap (PBS) operations, which is much more expensive and can dominate the overall computational cost.
Activation functions play a vital role in neural networks. Functions like sigmoid require complex mathematical operations even in clear text, including exponentials and divisions, which are computationally expensive. On the other hand, the ReLU (Rectified Linear Unit) activation function has simpler implementations involving threshold comparisons, enabling more efficient execution on various hardware architectures. Univariate non-linear activation functions can be evaluated in TFHE through the Programmable Bootstrap (PBS) operation, although it incurs significant computational costs.
As the Concrete Python library for TFHE only supports integers natively, quantization becomes necessary when implementing neural network architectures: inputs, weights, and activations all need to be projected from floating-point values onto the integers. Furthermore, since any deep neural network with non-trivial activation functions will require PBS, the maximum precision of the TFHE table look-up implementation will be a limiting factor. At the time of the experiments, this was $2^{7}=128$ different values, i.e., integer 7-bit precision [^1] Later versions support up to 16-bit precision.. Precision also directly impacts execution time as larger table lookups imply slower inference.
Method
@@
which is then used to form a weighted sum of values as output, H=SVH=S\,V, exploiting that each row of $S$ is normalized to sum to one. The FFN has two fully connected layers with weight matrix multiplication and ReLU activation,
@@
where $W_{1},W_{2}$ are weight matrices and $b_{1},b_{2}$ are bias vectors. We use the short-hand notation $x^{+}=\max(0,x)$ for the ReLU function. Each Transformer block typically also has one or more layer normalizations (Ba, Kiros, and Hinton 2016) that stabilize the gradient flow and contribute to regularizing the training.
The Inhibitor alternatively proposes to calculate attention scores as
@@
Results
[…]
stats_beta4_2_0_2_2.txt
Table 2: Parameters and circuit bit size set by Concrete TFHE compiler for four different sequence lengths (T) processed by the two alternative Transformer attention mechanisms. Note the difference in bit-precision that is required (last two columns).
Benchmark comparisons.
The adding problem was introduced by (Hochreiter and Schmidhuber 1997) to test long-term memory for RNNs. It takes two inputs: a sequence of random numbers and a two-hot sequence (which, in our experiments, each is 100 long). The ground truth is simply the dot-product of the two inputs as vectors, which, rather obviously, is not a challenging task for a conventional (dot-prod-based) Transformer. It is encouraging that the addition-based Inhibitor performs just as well on this task (which is hard for, e.g., simple RNNs).
We also train one-layer Transformers for the MNIST handwritten digit recognition task (LeCun and Cortes 2010) and IMDB movie-review sentiment analysis task (Maas et al. 2011), which are simple go-to benchmark task for image classification and text analysis, respectively. Although far from SotA, we note that our simple models achieve decent accuracy for both attention mechanisms, where the differences in performance are not significant.
The final task, labeled IAMW in Table 1, uses the IAM Handwriting Database (Marti and Bunke 2002), which is a collection of handwritten words written by more than 700 writers. The model first applies two convolutional layers followed by a Transformer and uses Connectionist Temporal Classification (CTC) loss (Graves et al. 2006) as an endpoint layer to predict the text. Edit distance (Navarro 2001) was used as the evaluation metric. Again, the differences are small (and not significant at 95% confidence).
@@ Results
For the plaintext experiment, we used integer 16-bit arithmetics implemented in the Rust programming language, which gives detailed low-level control over circuits and supports advanced time benchmarking through the Criterion package. The encrypted implementation uses the TFHE scheme (Chillotti et al. 2019) for Fully Homomorphic Encryption on the Torus as supported in the Concrete library for Python (Zama 2022). TFHE permits the evaluation of arbitrary single-variate activation functions under homomorphic encryption through the Programmable Bootstrap (Chillotti et al. 2015). We here used considerably smaller networks where the embedding dimension was limited to size 2. The encrypted circuits were implemented for integers with up to 8-bit precision.
@@
timing_short_beta4_2_0_2_2.txt
@@
The results indicate that the proposed Inhibitor mechanism can have a significant advantage, with i) 30%–50% saving for the plaintext implementation on CPU as per Table 3, and ii) a factor 3–6 under encryption with TFHE as per Table 4. Timing estimates are averaged over 20 repeated experiments and significant at the 95% confidence level. Table 2 displays compiler parameter settings for the differnt encrypted circuits. Note from the last two columns that the dot-prod based variant requires up to two bits higher precision than the Inhibitor. It also requires about twice as many PBS, which further contributes to it slower performance.
@@ Conclusions.
The proposed Inhibitor Transformer allows straightforward quantization by replacing conventional dot-product attention by a mechanism based on absolute value (eq. 5), subtraction and ReLU function (eq. 6) that all directly support an integer implementation. Our scaling experiments indicate that it has considerable potential for computational savings both for Transformers processing plaintext data as well as ciphertext. The speedup is obtained by replacing the query-to-key matrix multiplication and the subsequent Softmax with a light-weight Manhattan distance attention score and the inhibitor activation.
While experiment results are promising on simple training tasks, for future work, it is necessary to examine performance under more challenging settings, for example, by pre-training a much larger Transformer model and testing on modern NLP and image recognition benchmark tasks, both in plaintext and under homomorphic encryption.
@@
Signed Inhibitor
The inhibition mechanism described in equation 6 can only produce a non-negative attended value matrix $H$, while the conventional dot-product attention can select from $V$ that have both positive and negative entries. We can rewrite inhibition to handle signed values simply as
@@
Implemention
A naive implementation of equation 5 and 6 relies on expanding the dimension of the broadcasted differences, before passing the result through an $\mathrm{abs}$ or $\mathrm{ReLU}$ function and then summing over an inner dimension. This uses large amounts of memory and should be avoided.
@@
For the inhibition step of equation 6 we use a simple trick based on
@@
Relation to Prior Work
This version extends the original AAAI2024 submission (Brännvall 2024) by adding material focused on the implementation and evaluation of the Inhibitor attention mechanism under Fully Homomorphic Encryption (FHE), using the TFHE scheme on the Torus. In addition to the original benchmark results, this paper provides a more detailed cryptographic context, including parameter selection and performance under encryption. While the core methodology remains unchanged, the scope has been broadened to put more focus on privacy-preserving machine learning applications.